how to solve 3x + y - 2 = 0 and 2x - y - 3 = 0

Found 2 solutions by greenestamps, mahikab:
Answer by greenestamps(13200)



"3x + y = 2 = 0" is nonsense....

Re-post, taking the time to make sure you show the problem correctly.


Answer by mahikab(11)

We can use systems of equations!
We can keep the variables on the right side of the equation, while keeping the constants on the right:
So 3x + y - 2 = 0 becomes 3x + y = 2
So 2x - y - 3 = 0 becomes 2x - y = 3
Now we can use systems of equations:
3x + y = 2
+ (2x - y = 3)
-----------------
3x+2x + y-y = 2+3
5x + 0 = 5
x = 1
Since we know what one of the variables equal, we can plug in the value of x in one (or both) of our equations.
3(1)+y = 2
y = 2-3
y = -1
Let's check if it works for our other equation as well!
2(1)-y = 3
2-y=3
-y = 3-2
-y = 1
y = -1
So, the answer to our problem is:
x = 1
y = -1
Also could be written as (x,y) -> (1, -1)